OpenStudy (anonymous):

Evaluate 7m – 3p for m = 3 and p = -5

OpenStudy (anonymous):

All u need to do is multiply

OpenStudy (callisto):

Put m = 3 and p = -5 into 7m – 3p 7m – 3p = 7(3) – 3(-5) = 21-(-15) = 21+15 = 36

razor99 (razor99):

excellent answer

razor99 (razor99):

@Callisto

OpenStudy (anonymous):

7m-3p 7(3)-3(-5) 21-(-15) 21+15 36